Archived 2013 IT Notices

12/9/2013 - 9:33 AM

Recently we have received a rash of emails requesting that people enter their userID and password to manage their email space.  Such emails are commonly called phishing emails and are very well engineered messages designed to trick you into revealing your userID and password. Phishing emails escalate around the holidays. DO NOT RESPOND to such emails.  As soon as we become aware, we block them.  If we identify that someone’s account information has been revealed we disable the account. Responding to a phishing email can lead to identity and university data theft. It also causes other organizations to block email coming from LSUHSC.EDU which interferes with critical business at our university.

Here are two simple guidelines to keep from being tricked:

1)  LSUHSC-NO IT already has the necessary administrative privileges to manage the LSUHSC-NO IT infrastructure.  We don’t know your password and don’t need to know it to get our job done.  We would never ask for it, especially in response to an email.

2)  Never divulge your SSN, bank account or other such sensitive information via an email.  LSUHSC-NO already has all the information needed for administrative purposes.

Thanks to all who have picked up that an email was a phishing email and took the time to let us know about it.  If you ever receive a phishing or spam email and would like to report it, please forward it as an attachment to and we will investigate.

12/6/2013 - 10:02 PM

The Production PeopleSoft Campus Solutions application will be unavailable Saturday, Dec. 7th  from 7:30 am to 1:00 pm for maintenance.  During the maintenance, Academic Self-Service will be unavailable.  An email notification will be sent when the system is available.

11/27/2013 - 2:02 PM

The Production PeopleSoft HR/Payroll application will be unavailable Monday, Dec. 2nd  from 5:00 pm to 9:00 pm for maintenance.  During the maintenance, Employee Self-Service will be unavailable.  The HR/Payroll Reports database will be unaffected by this maintenance.   An email notification will be sent when the system is available.

08/08/2013 - 11:02 AM
Power has been restored to the Clinical Education Building at 1542 Tulane
Avenue, but it is still uncomfortable. Medical School faculty and staff should
continue working at their present locations. We will keep you updated.
07/26/2013 - 12:05 PM

We had 2 remaining issues, and believe that those are now resolved.  There was an issue logging into the CWI (Citrix Web Interface). That issue was resolved by rebooting the browser servers that were on the affected SAN.  

Our Moodle web servers were also not working properly, but that issue has also been resolved.  

Please contact the help desk, if you are experience any issues.  

07/26/2013 - 10:59 AM

We believe the issue has been resolved, and it was related to fiber connectivity between a SAN and a fiber switch.  Please report any new issues to the help desk.

7/26/2013 - 10:35 AM

We are currently experience response time issues with multiple applications. We are invesitgating the cause of the issue, and will udpate, as we get additional information.

06/26/2013 - 13:44

The process to correct the data in the GAL has completed, and all information should once again be correct.

06/26/2013 - 12:38 PM

A process is running to correct the GAL problems.  An additional update will be provided when the process completes.

06/25/2013 - 09:40

There is a problem with data in the Global Address List (GAL) this morning and incorrect information is contained in the GAL for many individuals.  We are actively investigating the issue and are working toward correcting the issue as quickly as possible.  Please pass this information along to the user community.

05/20/13 - 19:34

All services have been restored. The ATS for network switches failed and it did not switch over correctly when Venyu turned off power to Side A.

05/20/13 - 19:00

Due to power issues at Venyu, multiple HCSD applications are not working properly. Additional updates to follow.   Exchange is one of the applications affected.

/17/13 - 2:42 PM

At 3 am this Saturday, May 18th, we need to take one of PSFILESRVR’a volumes off-line, in order to complete maintenance that is required to clean up a minor issue that resulted from a Microsoft patch that was released in April.   (The patch in question has since been uninstalled.)   The only volume affected by this maintenance is PSFILESRVR’s “Z-drive”, which only impacts users when working inside of the Citrix Desktop or PS Desktop.    We expect the Z-Drive to be down for maintenance from approximately 3 am – 6 am Saturday.

During this maintenance, users will find that any personal shortcuts that they have saved on their personal “desktop” in Citrix will be missing, as well as any shortcuts that they have pinned to the taskbar or start menu inside of Citrix.  If they log on after 3 am, they will likely also receive a message about “Z:\Desktop” being ”unavailable”.     For most PS Desktop users, this may be the full extent of the impact.   However for the Citrix (clinical) “Desktop”, the impact will be somewhat greater.  Due to the way the TEMP variable in Desktop sessions is redirected to “Z:\Temp”, clinical Desktop users may encounter a variety of errors during the maintenance.   Therefore we recommend avoiding use of Desktop during the outage if possible.

5/9/13 - 10:15 PM

Maintenance has been completed on the server that hosts the V-drive.  If problems occur connecting to the V-drive, please log off and then back on.

5/9/13 - 3:47 PM

Tonight from 10:00 - 10:30 pm the file server that hosts the V-drive will be down for important maintenance.  During this time, applications such as PeopleSoft that depend on the V-drive will be unavailable.  In addition CLIQ users will need to access CLIQ from one of the methods other than running from the V-drive.   Citrix users can expect to receive error messages associated with the absence of the V-drive during the maintenance.  Additional messages will follow immediately before and after the maintenance.

04/30/13 - 09:58 AM

Following the completion of scans of the H:, I:, and J: drives, we are receiving reports from users whose home shares, O: drives, are inaccessible.  It appears that the CHKDSK process "repaired" security information on some set of files and folders on the drives that were scanned.  At this time we do not know for certain how many users are affected, but the current estimate is a few hundred based on reports submitted thus far, but we cannot be certain of this estimate.  We do not have a method to check permissions proactively and are handling the issues as they are reported.

04/30/13 - 1:10 AM

The CHKDSK of the H: drive on ENTFILESRVR has completed. The CHKDSK process did not attempt to scan any of the other system drives and thus the process did not have to be aborted. ENTFILESRVR is online and drives are mapping successfully. Work with continue with Microsoft in the hope of determining why the CHKDSK was needed on three of the server’s drives. This is the last planned update on this issue.

04/29/13 - 06:08 PM

The CHKDSK of the H: drive on ENTFILESRVR continues to run.  When the H: drive scan completes the CHKDSK will be aborted.  The current estimate for aborting the CHKDSK is between 12:00 a.m. and 3:00 a.m. with the expectation that the server will be online before the start of the workday on Tuesday.

04/29/13 - 01:22 PM

The CHKDSK process on ENTFILSRVR is inspecting the H: drive.  When the process completes on the H: drive the scan will be aborted.  Once the CHKDSK process is aborted the server should complete the boot process and the shares and files become available.  Current availability estimate is between 3:00 p.m. and 5:00 p.m.   We are waiting for the current drive to complete in the hope of avoiding a restore of the drive.

04/29/13 - 10:12 AM

The CHKDSK process on ENTFILESRVR continues to run.  We are working with Microsoft to develop a plan to bring the server online without allowing the CHKDSK process to complete.  At this time O: drives for LSUHSC-NO and ILH are unavailable.  Departmental T: drives are unavailable for many LSUHSC-NO users, but should be available for ILH.  Users may continue to experience problems with Outlook in Citrix.  Interrupting the CHKDSK process could result in corruption of the drive being scanned and require a restoration of data for that drive.  An additional update will be provided when we have completed work with Microsoft on a plan to bring the server online without CHKDSK running to completion

04/28/13 - 11:32 AM

A critical Enterprise server is having storage problems, and at this time, user's home shares on the O: drive, and department shares on the T: drive are unavailable. Also, users may experience problems using Outlook via Citrix. Alternative would be to use webmail at

03/21/13 - 04:04PM

Postmaster will perform maintenance on the New Orleans Mailbox servers during the maintenance interval between 3 AM and 6 AM this Saturday, March 23. The aggregate downtime will be less than one minute per mailbox as each server is failed over and then rebalanced.  An update will be sent when this maintenance is complete.  

03/20/13 - 07:45 AM

The LSU Health Remote Access Portal ( was updated around 3am this morning. After the update, users trying to access VPN using the Network Connect option, would get stuck in a loop trying to download hfnetchk6b.xml. The problem was resolved around 7:15am.

03/20/13 - 06:52 AM

Power has been restored to the CEB building at 1542 Tulane Avenue, and the building is now operational. The problem – a fault in a manhole at Julia and South Rampart streets – was located and repaired overnight, and normal activities have resumed. Thank you for your patience and cooperation during this process. 

03/19/13 - 10:58 AM

Clinical Education Building 1542 Tulane power outage expected to last another 24 hours.

3/19/13 - 09:07 AM

The Clinical Education Building in New Orleans is currently without power. Entergy is working on the problem and no timeframe has been given when the power will be restored.

03/15/13 - 09:50 AM

AT&T is reporting that the damaged cable adjacent to Sister Stanislaus Memorial Hall that affected the LSUHSC Foundation, Environmental Health & Safety, Wellness Center, Stanislaus Hall Housing Office, and all Stanislaus Hall elevators and emergency phones has been repaired. If your phone line was impacted, please verify that it has dial tone and report issues to .

03/14/13 - 5:23 PM

According to AT&T, an underground 1,200 pair cable close to Stanislaus Hall was damaged which has impacted telephone service to this building.  AT&T is working to repair the damage as quickly as possible.